Sol Summary – April 18th

Crew 280 Sol Summary Report 18-04-2023

Sol: 2

Summary title: "Hypatia on Air: Mars, Camera, Action"

Author’s name: Mariona Badenas-Agustí, Commander.

Mission status: Nominal

Sol activity summary: The crew woke up very early in order to welcome a media visit from Earth. Before this visit, some members had the opportunity to take their first shower, which was a nice reminder of the luxuries that we have on Earth and that we sometimes take for granted. After everyone got ready, we had breakfast and started working on our experiments.

At about 11:15AM, Mariona Badenas-Agusti (Crew Commander), Carla Conejo (Executive Officer), Neus Sabaté (Mission Specialist), and Núria Jar (Journalist) conducted an EVA. The original plan was to drive to Cowboy Corner, but we requested a change to the initial plan due to the high winds around the station. Our plan was approved, and the EVA team drove to Marble Ritual instead, using Curiosity and Spirit rovers.

The goal of today’s EVA was to take geological samples for Neus’ research experiments and get more familiar with the EVA protocols and spacesuits. At the same time, the EVA offered the media team a chance to obtain beautiful pictures and high-quality drone videos. We are looking forward to seeing their results.

At around 1PM, the EVA group returned to the Hab and had a delicious meal with the rest of the team. Laia Ribas (Greenhab Officer) and Cesca Cufí (Crew Engineer) prepared pasta with roasted vegetables and parmesan cheese and fruit. We then rested a bit before everyone continued to prepare their scientific experiments.

Soon after 5PM, the media team left the Hab and the crew participated in Carla’s psychological experiment. This evening, we will get in touch with Mission Support, have our second dinner together, and organize a debate about the future of Hypatia and our Association.

Look ahead plan: These past few days have been so intense that most crew members have not had the chance to finish the setup of their experiments. Now that our team is alone on Mars, everyone is excited to finally kick off their research without any interruptions.

Anomalies in work: Non-functional robotic telescope.

Weather: Sunny, but very high winds. Low 54F/High 71F.

Crew physical status: Everyone is feeling well, and our Crew Engineer is feeling much better than yesterday. She is nearly recovered!

EVA: We conducted an EVA to Marble Ritual in order to collect iron oxide samples.

Reports to be filed: Sol Summary, EVA request, EVA Report, HSO Report, Operations Report, Greenhab Report, Journalist Report, Astronomy Report, Pictures of the Day (with Photo of the Day)

Support Requested: We need new batteries for the gas detectors. We would also like to notify that all the stove knobs are broken and constitute a safety hazard (see HSO and Operations Report).

Operations Report – April 18th

Crew 280 Operations Report 18-04-2023
SOL: 2
Name of person filing report: Neus Sabaté – Mission Specialist (Assisted by Cesca Cufí – Crew Engineering)
Non-nominal systems: robotic observatory
Notes on non-nominal systems: Replacement parts are scheduled to be installed during crew 261 rotation.
Spirit rover used: yes
Hours: 216.7
Beginning charge: 100%
Ending charge: 88%
Currently charging: no
Opportunity rover used: no
Hours: 120.7
Beginning charge: –
Ending charge: –
Currently charging: no
Curiosity rover used: yes
Hours: 229.8
Beginning charge: 100%
Ending charge: 92%
Currently charging: no
Perseverance rover used: no
Hours: 261.1
Beginning charge: –
Ending charge: –
Currently charging: no
General notes on rovers: non
Summary of Hab operations:
Water use: can not be estimated because water tank has been partially refilled
Water (static tank): 280 gallons
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: yes
Summary of internet: used from 18:00 to 21:00 for personal communication and reports.
Summary of suits and radios:
Suits: spacesuits 1,2,5 and 6 have been used during the EVA. All nominal.
Radios: all operative during EVA
All radio and spacesuit batteries checked before EVA: ok.
Radio batteries charged after EVA.
Spacesuit batteries plugged.
Summary of GreenHab operations:
Water use: 14 gallons
Heater: off
Supplemental light: off
Harvest: 191 gr of cucumber
Radios were found discharged: the jack plug of charger was half disconnected so the radios were not properly charged. This has already been solved.
Summary of Science Dome operations:
Laia Ribas (Green Hab Officer) prepared the setups for her experiments.
Neus Sabaté (Mission Specialist) brought rock samples from EVA and prepared the setup for further analysis of electroactive iron contents.
Dual split: off
Summary of RAM operations:
Cesca Cufí (Crew Engineer) used the RAM to start mounting her experiment setup (homemade biological safety cabinet).
Summary of any observatory issues: none
Robotic observatory: Not working. Replacement parts are scheduled to be installed during crew 261 rotation.
Summary of health and safety issues:
Gas stove knobs dammaged: the gas stove knobs are damaged and they don’t fit properly the valves axis, resulting in a serious hazard. Because of this reason, there was a gas leak during cooking lunch, while during the EVA. The two cooks detect immediately the leak but could not stop it immediately as the knob was not working and the other knobs neither. The leakage was stopped several minutes after by the crew engineer (cooking at this time) that turned off the knob axis using a pliers that she had in the Hab, while one crew member went to the RAM to pick a gas sensor to monitor the concentration of gas. After the EVA, to ensure the safety, the Crew Engineer explained the issue to all crew members. The HSO has recommended limiting the use of the stove to the single burner that has an actual functional knob while we find a way to repair the rest. A gas detector has been placed in the kitchen to measure the gas concentration in case we need to use several burners, to ensure all burners are off when finishing cooking. All crew members are aware and know the procedure to follow.
Questions, concerns and requests to Mission Support:
The gas detector that has been used to detect the gas leak from the kitchen stove has a very low battery level. Also the second one we found in RAM. We have looked for replacement but could not find AAA batteries in the RAM.
In case Mission Control refuses the crew repairing/substituting the knobs, we require spare parts to replace all of them (x4). We can provide the knob measures and pictures.

Journalist Report – April 18th

Crew 280 Journalist Report 18-04-2023

Author: Núria Jar, Crew Journalist

While food, sleep, and extravehicular activity (EVA) protocols are now becoming the new routine for the Hypatia I crew members, they are still adjusting to the new living conditions. This morning, some crew members took a quick shower, a privilege on Mars, where water is a scarce resource on the Red Planet.

As the Hypatia I team is settling into its new home, some crew members have already started to set up their experiments to conduct high-quality space research in the Martian simulation environment. Their discoveries will contribute to the understanding of what it would be like to live and work on Mars.

The diverse backgrounds and expertise of the Hypatia I crew can allow them to approach research from multiple angles, which is important in making progress in different areas related to space exploration. By focusing on a diverse range of research fields such as astronomy, space biology, engineering, and outreach, they can gain a comprehensive understanding of the challenges and opportunities involved in exploring new planets.

The strong wind had a significant impact on the daily activities of the Hypatia I crew, requiring readjustments. Despite the wind speed being above the threshold of 20 miles per hour, four crew members still conducted an EVA, but they had to do so closer to the base than initially planned.

The decision to embark on an exploration was made because the documentary crew B23 arrived to the Mars Desert Research Station (MDRS) to film the crew’s activities and document their mission to Mars. The independent film company was on shooting while the Hypatia I crew continued with their simulation in the Martian environment. Two members of the crew are scientists from the Spanish National Research Council (CSIC), which has provided support for the documentary project.

Although it may sound strange from Earth, it is crucial for the crew of Hypatia I to respect the simulation they are in, in order to avoid any disruptions or damages to the experiment. This is especially important as they have already spent three days on their mission, and any deviation from the simulation could compromise the integrity of their research.

Crew members are indeed feeling exhausted, but they recognize the importance of prioritizing self-care and rest in order to ensure they can continue to conduct their mission safely and effectively. Despite the challenges, they find the experience of pushing the frontiers of science together stimulating and rewarding, and they remain committed to achieving their goals on this mission.

GreenHab Report – April 18th

Crew 280 GreenHab Report 18-04-2023

GreenHab Officer: Laia Ribas

Environmental control: Open during the day and heater during the night

Temperatures: 81.4 F

Hours of supplemental light: N/A

Daily water usage for crops: 14 gallons

Daily water usage for research and/or other purposes: 0.25 gallons

Water in Blue Tank: 169 gallons

Time(s) of watering for crops: 08:30, 17:00 h

Changes to crops: seed tomatoes are cultured

General watering of plants and crops was done today. There was high wind therefore the door was closed almost the whole day. Two experiments have started.

1- Tomatosphere
Principal investigator (PI) researcher: Ari Farrés
Supervision: Laia Ribas
Procedure: Two sets of tomato seeds have been planted. The first set (control) has always been on Earth, and the second set has experienced different radiation exposure, as they have spent some time in outer space. The goal of the experiment is to compare the growth of the two sets of seeds and see if we are able to determine which ones have been in space. We currently do not know which of the two sets has been in space.

2. Seeding for the crew
Principal investigator (PI) researcher: Laia Ribas
Procedure: The aim of the project is to provide fresh food to the crew. A total of 75 g of three bean Mix Nature Jim’s Sprouts have been cultured. For 5 hours the seeds were watered. Following that, the water was discarded and the seeds were cultured in wet tissue paper in a plastic box half covered.

Harvest: 190 g cucumber

Support/supplies needed:
I would like to get instructions on what is the best procedure to follow if there is a windy day with high temperatures as the door can not be properly subjected within the current format.
I noticed that the two thermometers of the Greenhab are not in concordance.

EVA Report – April 18th

Crew 280 EVA Report 18-04-2023

EVA #3

Author: Mariona Badenas (Crew Commander and Astronomer) and Carla Conejo (Executive Officer & Biologist)

Purpose of EVA: The main goal of the EVA was to collect iron oxide samples for Neus Sebaté’s (Mission Specialist) research project, which consists of building Martian batteries with iron materials. During this EVA, Carla Conejo (Executive Officer & Biologist) supported Neus in the sampling process, Mariona Badenas (Commander) was in charge of the navigation, and Núria Jar (Journalist) documented the EVA for her podcast. The EVA group was accompanied by two media teams (documentary and Catalan TV press teams).

Start time: 11:20AM

End time: 12:47PM

Narrative: The EVA was planned for 11AM, but it finally started at 11:20AM. It took 30 minutes for the crew to get ready, and during this time, both the media and the documentary team recorded the preparation of the EVA. After performing 5 minutes of depressurization in the front airlock, we drove Curiosity and Spirit to Marble Ritual to gather iron-rich rock samples for Neus Sabaté (Mission Specialist). With these samples, Neus will try to build batteries with Martian-like materials.

After a smooth but windy drive to our destination, we parked the rovers and walked east of Marble Ritual, where we located many black and reddish rocks with iron oxide materials. Mariona Badenas (Commander & Astronomer) and Carla Conejo (Executive Officer & Biologist) assisted Neus with the rock sampling process, while Núria Jar (Journalist) reported the whole process with beautiful pictures (see “Pictures of the Day”).

During this time, the broadcasting team filmed the crew, although all EVA teams made the effort to disregard their presence to respect the simulation. We completed the scientific goals of the EVA at 12:15PM, after walking and sampling for 1 hour at about 400m from where we parked the rovers. After that moment, we communicated our intended route back to the rovers to the media team through a radio that we provided them, and they followed us with a drone.

We then drove back and returned to the Hab, where we followed the 5-minutes pressurization protocol. The first media team abandoned the area soon after, while the documentary team entered the base a few minutes after the EVA team returned to the base. Although we had a windy expedition, we experienced no problems during this EVA.

Destination: Originally Cowboy Corner, but Marble Ritual.

Coordinates (use UTM WSG 84): 519000 E, 4250500 N

EVA participants: 4: Mariona Badenas (Commander) (EVA Leader), Carla Conejo (Executive officer & Biologist), Neus Sabaté (Engineer), Núria Jar (Journalist).

Road(s) and routes per MDRS Map: Cow Dung Road 0110.

Mode of travel: Drive on Cow Dung Road 0110 to Cowboy Corner then walking around no farther than 400 meters for sampling.

Astronomy Report – April 18th



Crew 280 Astronomy Report 18-04-2023

Crew 280 Astronomy Report 18-04-2023
Name: Mariona Badenas
Crew: 280 (Hypatia I)
Date: 18-04-2023

Robotic Telescope Requested (choose one) MDRS-14 MDRS-WF MLC
Objects to be Imaged this Evening: Planetary Nebula NGC6853 (astrophotography filters, x5 exposures of 300s each)
Images submitted with this report: None

Problems Encountered: None


Solar Features Observed: N/A
Images submitted with this report: N/A
Problems Encountered: N/A

Copyright © The Mars Society. All rights reserved. | Main Site